An open flask sitting in a lab fridge looks empty, but we know that actually it is filled with a mixture of gases called air. If the flask volume is 3.50 L, and the air is at standard temperature and pressure, how many gaseous molecules does the flask contain?
1
Expert's answer
2018-04-07T08:00:10-0400
Since the flask with air is in normal conditions, we can apply Avogadro's law, according to which equal amounts of different gases, taken at the same temperatures and pressures, contain the same number of molecules.
We also know the constant Avogadro - the number of molecules in 1 mole of matter and it is equal NA6,02x1023.
Also, therefore, from Avogadro's law, 1 mol (equal number of moles) of any gas for the same - isobaric and isothermal - conditions occupies the same volume. In particular, under normal conditions, that is, at 0 ° C (273 K) and 101.3 kPa, the volume of 1 mole of gas is equal to VM=22.4 liters.
Using these well-known theoretical data, one can proceed to calculations:
N=v×N_A=V/V_M ×N_A=3.5/22.4×6.02*〖10〗^23=9.4063*〖10〗^22

Answer: the flask contains 9.4063*〖10〗^(22 ) molecules of air